# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

### v1

@@ Abstract.
This study explores the nuanced capabilities and inherent biases of Recommender Systems using Large Language Models (RecLLMs), with a focus on ChatGPT-based systems. It studies into the contrasting behaviors of generative models and traditional collaborative filtering models in movie recommendations. The research primarily investigates prompt design strategies and their impact on various aspects of recommendation quality, including accuracy, provider fairness, diversity, stability, genre dominance, and temporal freshness (recency).
Our experimental analysis reveals that the introduction of specific system roles and prompt strategies in RecLLMs significantly influences their performance. For instance, role-based prompts enhance fairness and diversity in recommendations, mitigating popularity bias. We find that while GPT-based models do not always match the performance of CF baselines, they exhibit a unique tendency to recommend newer and more diverse movie genres. Notably, GPT-based models tend to recommend more recent films, particularly those released post-2000, and show a preference for genres like ‘Drama’ and ‘Comedy,’ and ‘Romance’ (compared to CF ‘Action’, ‘Adventure’) presumably due to the RecLLMs’ training on varied data sets, which allows them to capture recent trends and discussions more effectively than CF models.
Interestingly, our results demonstrate that the Simple and Chain of Thought (COT) paradigms yield the highest accuracy. These findings imply the potential of combining these strategies with scenarios that favor more recent content, thereby offering a more balanced and up-to-date recommendation experience. This study contributes significantly to the understanding of emerging RecLLMs, particularly in the context of harms and biases within these systems. It highlights the complexities and potential of prompt design in impacting recommendation outcomes and paves the way for future research in optimizing accuracy, fairness, and user experience in such systems.
@@ 1. Introduction
Recommender systems have increasingly become integral in various domains, offering substantial benefits for stakeholders in two-sided, or broadly, multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019). These systems alleviate user information overload and enhance producer sales and revenue. From a learning viewpoint, recommendation models can generally be classified into two categories: ‘discriminative’ and ‘generative’ models (Deldjoo et al., 2022).
Discriminative models, often considered the traditional approach in recommendation systems, function by categorizing and differentiating to inform decisions about user preferences. For example, in the movie domain, these models can be compared to expert critics that are adept at discerning the distinctive characteristics that separate one type of item (movie) from another, given the audience. In other words, this differentiation is informed by factors such as user historical ratings, genre preferences, and other relevant attribute of items. Mainstream collaborative filtering (CF) recommendation models, such as matrix factorization (MF)-based models and more complex neural variations such as NeuMF (He et al., 2017), operate on this principle. Similarly, graph-based models, such as NGCF (Wang et al., 2019) and LightGCN (He et al., 2020), leverage the relational structure of the data to improve the quality of recommendations. Regardless of the methodological approach, traditional discriminative models are heavily based on knowledge derived from limited datasets and specific domains to calculate recommendations and inform decisions.
Generative models represent a transformative approach, focusing on understanding the underlying structure and distribution of data to generate authentic-looking content. Their application in recommender systems (RS) exemplifies a shift from simple preference matching to a more nuanced understanding of user behavior and interests. In the context of movie recommender systems, these models can be likened to hypothetical curators who not only distinguish between various items but also comprehend the overarching patterns and tastes of the audience. Thus, these models go beyond mere differentiation; they synthesize and anticipate user preferences, providing tailored recommendations.
Large Language Models (LLMs), exemplified by the Generative Pretrained Transformers (GPT), recently heralded a new era in Natural Language Processing (NLP). These Transformer-based models, such as BERT, GPT series, PaLM series, etc., are characterized by their vast number of parameters and are trained on extensive datasets using self/semi-supervised learning techniques. Their training on a wide range of internet text data endows LLMs with the exceptional ability to generate contextually relevant and coherent text in response to input prompts. Different from traditional recommendation systems, LLM-based models excel in capturing contextual information, comprehending user queries, item descriptions, and other textual data more effectively (Geng et al., 2022). Consequently, specifically-designed LLMs for recommender systems (RecLLMs) can harness LLMs to extract high-quality representations of textual features and leverage extensive external knowledge encoded within LLMs (Liu et al., 2023) to offer more nuanced insights into understanding user preference and profile. This is a significant leap forward compared to traditional RS. This aligns well with the overarching goal of delivering personalized, context-aware recommendations that reflect individual tastes and behaviors.
In general, LLMs can contribute to recommender systems in three principal ways (Wu et al., 2023; Li et al., 2023b; Xu et al., 2024). First, LLMs serve as direct recommenders by generating personalized content and suggestions, stemming from their advanced understanding of context and user queries (Geng et al., 2022; Wu et al., 2023). Second, they enhance traditional models by extracting rich semantic representations from textual data. This aspect is particularly important in addressing the data sparsity issue in historical interactions (Da’u and Salim, 2020). Third, LLMs act as complex simulators in the recommendation process, effectively testing and refining the system’s predictions (Liu et al., 2023). The current work in hand primarily concentrates on the first application, that is using LLMs as recommender, with ChatGPT serving as the case study.
The primary aim of this research is to enhance our understanding of “Provider Fairness within ChatGPT-based RecLLMs”. We focus on uncovering and understanding a range of biases – mostly underexplored in existing literature –, and examining how these biases manifest in comparison to traditional Collaborative Filtering (CF) models. This study looks into the dynamics of these biases, providing a detailed contrast between the two types of systems, and thereby contributing to a more nuanced understanding of the complexities inherent in modern RS.
In particular, this research diligently identifies and analyzes inherent biases inherent in these two types of systems, such as (i) the preference for popular movies, (ii) the balance between old and recent movie recommendations, (iii) the tendency to recommend certain movie genres, (iv) the consistency/stability of recommendations over time. For instance, GPT-based models were found to have a penchant for suggesting recent films like those released post-2000, exemplified by their frequent recommendations of newer titles such as ”Inception” (2010) and ”The Dark Knight” (2008), compared to CF models that often favored classics like ”The Godfather” (1972) and ”Casablanca” (1942). Furthermore, genre dominance was evident, with GPT models showing a distinct inclination towards ‘Drama’ and ‘Comedy’, whereas CF models predominantly recommended ‘Action’ and ‘Sci-Fi’ films. Another crucial aspect of this research is exploring the impact of prompt engineering in a zero-shot learning setting on the performance of RecLLMs in terms of personalization, diversity, recency, and fairness. Zero-shot learning refers to the model’s capacity to handle tasks for which it has not been explicitly trained, relying instead on its general language understanding and adaptability.
Here our study seeks to provide an answer to the following question: Can prompts serve as a controllability tool to steer the recommendation process in a specific direction (increase accuracy, reduce provider unfairness) based on the interests of providers or other stakeholders, and what potential harms are associated with this approach?
@@
Figure 1. Conceptual idea behind our work
@@
Example. We illustrate the impact of different “prompt design” strategies on the output of ChatGPT-based recommendation systems through a clear example, as shown in Figure 1. We use a randomly chosen user profile from the MovieLens dataset to demonstrate this effect. The user’s historical interactions with the system include a diverse range of movies such as “Toy Story” (categorized with genres Adventure—Animation—Children—Comedy—Fantasy), “If Lucy Fell” (Comedy—Romance), and “Hard Target” (Action—Adventure—Crime—Thriller).
@@
To highlight the differences in recommendations based on the design of the system, we designed a total of seven prompts categorized into three distinct classes. These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the system’s outputs can vary significantly depending on the design strategy employed.
@@
Promoting fairness in RecLLM: To combat popularity bias, we tested two approaches: (1) Instructional prompts: Statements like “Focus on fair recommendations, balancing popular and lesser-known movies” were integrated into various scenarios, aiming to influence RecLLM’s output. (2) Role-based directives: We explored directly incorporating fairness into the “system’s role”, testing its effectiveness compared to solely instructional prompts. This experiment seeks to identify the most effective strategies for achieving fair recommendations in RecLLM. In sum, our work demonstrates how strategically designing a prompt’s semantic intent and structure can impact a recommender system’s performance across accuracy, provider fairness, diversity, and temporal freshness (recency).
Contributions. The following is a list of proposed contributions in the current work:
  1. (1)
@@
    Enhanced Analysis of Prompt Design in Zero-Shot RecLMMs. This research provides an in-depth study into how the design of prompts (in terms of intent, structure, and system role) affects the performance of ChatGPT-based recommendation systems. It investigates how different prompt design impacts factors such as accuracy, fairness, diversity, and temporal freshness in recommendations. This study involves experimental designs that compare various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and analyze their impacts on the recommendation of niche versus popular content, as well as the balance between old and recent movies.
  2. (2)
@@
    Identification and Analysis of Different Biases in Zero-Shot RecLLMs. This work identifies and analyzes numerous biases that may be present in RecLLMs such as ChatGPT, thus laying the groundwork for investigating bias mitigation strategies. Key biases identified under scrutiny include popularity bias, genre preference bias, and temporal (recency) bias. We study how these biases uniquely manifest in LLM-based systems versus traditional Collaborative Filtering (CF) models.
  3. (3)
    Stability Analysis in Zero-Shot RecLLMs: This world also studies the stability of recommendations provided by LLM-based systems over time. Stability in this context refers to the consistency of recommendations given similar input conditions at different times. Our study analyzes and discusses factors that might impact stability (besides the parameters controlled by LLMs), such as changes in underlying data, model updates, and prompt variations. It would also explore the implications of stability for user satisfaction and trust in the system in real-life deployments.
@@
The structure of the paper is organized as follows. Section 2 presents the related work, focusing on the exploration of fairness in recommendation systems (RS) and pre-trained language models (LMs), as well as their applications in enhancing RS. Section 3 is dedicated to presenting a thorough evaluation of ChatGPT. This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1), the repeated experiments for stability analysis (Section 3.2), an understanding of the system’s role in ChatGPT-based RecLLM (Section 3.3), and the promotion of fairness (Section 3.4). Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.5). Following this, Section 4 outlines the experimental setup employed in our study. The results and key findings of these experiments are thoroughly discussed in Section 5. The paper concludes with Section 6, where we summarize our findings and outline potential avenues for future research.
@@ 2.1. Fairness in Recommender Systems
Fairness has emerged as a pivotal topic within the AI community in the last few years, arguably becoming the most scrutinized aspect across various branches of trustworthy AI, including fairness, security, privacy, and explainability. In RS, fairness has garnered significant attention due to the multi-stakeholder nature of these systems (Deldjoo et al., 2023; Ekstrand et al., 2019). Unfairness, even in its minimal form, can adversely impact various stakeholders, including consumers, producers, system designers, supply chains, and even the environment – the latter are often referred to as ‘side-stakeholder’.
@@
The body of literature on fairness in RS is diverse, encompassing multiple perspectives and dimensions. Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), have introduced various taxonomies to categorize these diverse aspects into several orthogonal or partially-orthogonal dimensions. Key dimensions recognized in the literature include the main stakeholder in question (e.g., consumer vs. producer), the target benefits associated with each (such as effectiveness vs. item exposure), the granularity of sensitive groups for assessing fairness (individual vs. group level), and other dimensions including the core definition of fairness, temporal aspects, and more. These dimensions offer a comprehensive view of the fairness landscape in RS, as detailed in these surveys.
@@
In the context of the present study, and with regard to positioning this work within the literature of fair Recommender Systems (fair-RS), we have introduced Table 1. This table is designed to categorize existing literature along two principal dimensions: the stakeholder in question (consumer vs. producer) and the nature of the core RS under scrutiny (traditional RS vs. those based on recent advancements in LLMs). We briefly review these dimensions in the following.
@@
Core RS under scrutiny. We distinguish between ‘traditional’ RS and those enhanced by ‘RecLLM’. This distinction is crucial because, while RecLLM, such as those based on GPT-like architectures, promise to significantly advance RS landscape with more nuanced and personalized recommendations, they also raise concerns about inherent biases. The vast and unregulated nature of internet data used for training LLMs raises concerns about inherent biases against specific races, genders, popular brands, and other sensitive attributes. For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands. Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways. Therefore, accurately measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies, ensuring that advancements in RecLLM do not amplify existing inequalities. This is the key goal of the current work in hand.
@@
  • Traditional RS. This column lists studies that have focused on traditional methods of recommender systems. They primarily rely on CF algorithms (possibly using side information of users and items) without the advanced natural language processing capabilities of LLMs. Within these works, some are dedicated to building evaluation frameworks for evaluating RS unfairness, while others focus on developing various mitigation strategies.
@@
  • RecLLM: This column spotlights the burgeoning research domain that merges language models (LMs) with RS representing a significant shift towards harnessing advanced NLP techniques to enhance the accuracy and relevance of recommendations. These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023a; Zhang et al., 2023). Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023a; Zhang et al., 2023), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
@@
Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023)
@@
Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023a)
@@
Stakeholder. As mentioned earlier, a dominant aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides. Within each of these segments, as you can observe, we can further categorize the literature based on which sensitive attribute groups are defined. For example, sensitive attributes such as the demographics of consumers (age, gender) and producers, as well as the popularity of items (on the produce side) are quite common focal points.
  • Consumer Fairness. This category is subdivided into various attributes like Activity, Demographics, Merits, and Others. It includes studies that focus on ensuring fairness among consumers of the recommender system based on these attributes. For instance, ensuring that recommendations are not biased towards a particular demographic group.
@@
The study by Zhang et al. (2023) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness. Their work introduces a novel benchmark, FaiRLLM, for evaluating the fairness of RecLLM, highlighting ChatGPT’s biases towards certain sensitive user attributes in music and movie recommendations. Conversely, the work by Li et al. (2023a) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation. This study investigates ChatGPT’s performance in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection. Similarly, research such as that by Di Palma et al. (2023) examines the personalization and popularity bias in RecLLMs, particularly in the context of ChatGPT and other LLMs. However, these studies do not address the intricacies of prompt engineering in RecLLMs, nor do they comprehensively address the various forms of biases studies in the current study. Our research goes beyond examining provider fairness based on popularity bias and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.

### v3

@@ Abstract.
This paper explores the biases inherent in ChatGPT-based recommender systems, focusing on provider fairness (item-side fairness). Through extensive experiments and over a thousand API calls, we investigate the impact of prompt design strategies—including structure, system role, and intent—on evaluation metrics such as provider fairness, catalog coverage, temporal stability, and recency. The first experiment examines these strategies in classical top-K recommendations, while the second evaluates sequential in-context learning (ICL).
In the first experiment, we assess seven distinct prompt scenarios on top-K recommendation accuracy and fairness. Accuracy-oriented prompts, like Simple and Chain-of-Thought (COT), outperform diversification prompts, which, despite enhancing temporal freshness, reduce accuracy by up to 50%. Embedding fairness into system roles, such as “act as a fair recommender,” proved more effective than fairness directives within prompts. We also found that diversification prompts led to recommending newer movies, offering broader genre distribution compared to traditional collaborative filtering (CF) models. The system showed high consistency across multiple runs.
The second experiment explores sequential ICL, comparing zero-shot and few-shot learning scenarios. Results indicate that including user demographic information in prompts affects model biases and stereotypes. However, ICL did not consistently improve item fairness and catalog coverage over zero-shot learning. Zero-shot learning achieved higher NDCG and coverage, while ICL-2 showed slight improvements in hit rate (HR) when age-group context was included. Overall, our study provides insights into biases of RecLLMs, particularly in provider fairness and catalog coverage. By examining prompt design, learning strategies, and system roles, we highlight the potential and challenges of integrating large language models into recommendation systems, paving the way for future research. Further details can be found at https://github.com/yasdel/Benchmark_RecLLM_Fairness.
@@ 1. Introduction
Context. Recommender systems are integral to various large-scale internet services, benefiting consumers, producers, and other stakeholders in multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019). The advent of generative models, particularly large language models (LLMs), holds the promise of offering better personalization experiences. LLMs enable conversational natural language (NL) interactions (He et al., 2023; Biancofiore et al., 2024), unlocking rich NL data like item descriptions, reviews, and queries within recommender systems (RS). Harnessing the general reasoning abilities of pretrained LLMs allows for addressing diverse, nuanced NL user preferences and feedback through highly personalized interactions. This contrasts with rigid ID-based methods that heavily rely on non-textual data. (Deldjoo et al., 2024a, b). This paper focuses on the emerging role of LLMs, specifically ChatGPT, in recommender systems and scrutinizes their biases, with a particular emphasis on “item-side fairness.”
Item-side fairness is important to ensure diverse item groups receive fair exposure, benefiting item producers such as micro-businesses in job recommendations and promoting content related to vulnerable populations. Recent research shows that LLM-based recommender systems (RecLLMs), with their reliance on semantic clues and generative capabilities, can introduce unique biases not present in conventional systems (Deldjoo and Di Noia, 2024; Jiang et al., 2024; Deldjoo et al., 2023; Deldjoo, 2024; Zhang et al., 2023a). Therefore, it is essential to investigate how these biases manifest and impact various stakeholders.
1.1. Background and Motivation
At a high level, the idea of generative modeling involves creating applications that not only make decisions based on data but also generate new data by learning patterns. This powerful idea has enabled various applications in AI disciplines, such as image generation, text synthesis, and music composition (Min et al., 2023; Koh et al., 2024; Ferreira et al., 2023; Zhang et al., 2023b). Generative modeling has recently gained prominence due to advances in model paradigms such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational Autoencoders (VAEs) (Kingma and Welling, 2013), Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020), and Transformer-based architectures such as GPT (Wei et al., 2022a; Bubeck et al., 2023) and other LLMs. In the recommender system (RS) field, generative models are not entirely new and have been used in various capacities for tasks such as data synthesis and augmentation (Chae et al., 2019; Yuan et al., 2020), model regularization (Liang et al., 2018; Sachdeva et al., 2019; Ma et al., 2019), and generating complex recommendation structures (Jiang et al., 2018; Liu et al., 2021). These advancements fall under Deep Generative Models (DGMs), which combine traditional probabilistic models with Deep Neural Networks (DNNs). The core strength of DGMs lies in their ability to model and sample from the data distribution they are trained on for various inferential purposes.
@@ 1.1. Background and Motivation
Figure 1. Conceptual idea behind experiment 1, prompt-design scenarios.
However, these models have experienced a resurgence of interest in the RS community, similar to trends in the broader AI landscape. This resurgence can be partly attributed to the introduction of LLMs such as ChatGPT, which have significantly improved the natural language understanding capabilities of these systems. LLMs demonstrate impressive in-context learning (ICL) and few-shot generalization abilities, making them valuable assets in both direct and indirect applications within recommender systems, where our goal here is to use them as direct recommenders (shown in Figure 1).
Notwithstanding their great success, the vast and unregulated nature of the internet data used to train LLMs raises concerns about possible biases against specific races, genders, popular brands, and other sensitive attributes that could be encoded in these networks. For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands. Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways. Another example could be the over-representation of content from certain geographical regions, leading to a lack of diversity in recommended media or news articles. Hence, the unchecked employment of these systems in commercial recommender systems may lead to unfair treatment of minority groups, reinforcing existing stereotypes, or exacerbating economic disparities.
This research explores different biases of ChatGPT-based RecLLMs, specifically on the provider side, to understand how these biases manifest in recommendations and affect various stakeholders. The primary goal of this research is two-fold:
  1. (1)
    To enhance our understanding of the general performance of ChatGPT-based RecLLMs in terms of accuracy, provider fairness, and other nuanced evaluation dimensions (genre dominance, temporal stability, temporal freshness); and
  2. (2)
    To determine whether factors such as prompt design and ICL type can be leveraged in RecLLM design to enhance item fairness without compromising accuracy;
To the best of our knowledge, the factors considered in this study are novel and have not been given enough attention in previous research. We believe that combining these studies into a single research effort opens up new avenues for future research in RecLLMs. Through two separate but extensive experiments, we introduce the following catalysts in the current work in hand:
  • Prompt Design. We examine how different prompt structures and instructions affect recommendation outcomes. As shown in Figure 1 and Table 3, we designed seven prompt scenarios classified into three categories: accuracy-oriented, beyond-accuracy, and reasoning. For instance, prompts instructing the model to diversify resulted in recommendations that were temporally fresher but not significantly novel or diverse, leading to a significant reduction in accuracy (e.g., NDCG) by 50%. In contrast, accuracy-oriented and reasoning-focused prompts, such as Simple and Chain-of-Thought (COT), performed much better.
  • System Roles. We further investigated the impact of assigning specific system roles, such as “acting as a fair recommender.” We observed that incorporating statements about fairness tends to improve system quality more effectively than including fairness directives directly within the prompt.
  • Stability Over Time. We considered that external factors such as trending data or algorithmic updates might introduce randomness into the responses of GPT-based models. To study this, each experiment was conducted five times specifically for GPT-based models. We assessed the consistency of recommendations over multiple runs to ensure reliability.
  • In-Context Learning (ICL). We analyzed the effect of zero-shot versus few-shot in-context learning (ICL) on provider fairness, focusing on demographic information impact on recommendations. We noted that incorporating user demographic data in prompts affected model biases and stereotypes. We structured the user profile into context, example, and demographic information parts. Our study highlights the nuanced effects of user profile attributes and sampling strategies on recommender systems.
@@
Example related to Experiment 1. We illustrate the impact of different “prompt design” strategies on the output of ChatGPT-based recommendation systems through a clear example, as shown in Figure 1. We use a randomly chosen user profile from the MovieLens dataset to demonstrate this effect. The user’s historical interactions with the system include a diverse range of movies such as “Toy Story” (categorized with genres Adventure—Animation—Children—Comedy—Fantasy), “If Lucy Fell” (Comedy—Romance), and “Hard Target” (Action—Adventure—Crime—Thriller).
@@
To highlight the differences in recommendations based on the design of the system, we designed a total of seven prompts categorized into three distinct classes. These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the outputs of the system can vary significantly depending on the design strategy employed.
@@
1.2. Contributions.
We designed two carefully-structured experiments focusing on classical top-k recommendation and sequential in-context learning. These experiments involved thousands of API calls to gather the findings presented in this work. The first experiment investigates how different prompt design strategies impact recommendation accuracy, provider fairness, as well as genre dominance and temporal freshness in classical top-k recommendations. The second experiment evaluates the effectiveness and fairness of RecLLMs in sequential in-context learning tasks, comparing zero-shot and few-shot learning scenarios across different datasets. This work makes several contributions to the field of recommender systems, as listed below:
@@
  • Enhanced Analysis of Prompt Design in Zero-Shot RecLLMs: We provide an in-depth study of how different prompt designs (in terms of intent, structure, and system role) affect the performance of ChatGPT-based recommendation systems. This includes an analysis of various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and their impacts on recommendation personalization, diversity, recency, and fairness.
@@
  • Identification and Analysis of Biases in Zero-Shot RecLLMs: Our research identifies and analyzes numerous biases present in RecLLMs, such as item fairness, genre preference bias, and temporal (recency) bias. We compare these biases with traditional CF models, providing insights into how they uniquely manifest in LLM-based systems;
  • Stability Analysis in Zero-Shot RecLLMs: We study the stability and consistency of recommendations provided by LLM-based systems over time. Specifically, we examine variations caused by changes in underlying data, model updates, and their implications on the quality of recommendations obtained by RecLLMs;
  • Generalization of Fairness Principles in Sequential Recommendations: We evaluate the effectiveness of RecLLMs in sequential recommendation tasks across different datasets, focusing on zero-shot versus few-shot learning. This analysis provides insights into the adaptability and robustness of these models in maintaining fairness across varied recommendation contexts.
  • Impact of User Profile Attributes and Demographic Information on Fairness and Accuracy: We explore the influence of user profile attributes (historical interaction sampling) and the inclusion of demographic information (e.g., gender, age-group) on recommendation fairness and accuracy, providing new insights into how revealing such information can affect model biases and stereotypes.
These contributions collectively enhance our understanding of prompt-based recommendation systems using LLMs, offering new perspectives on their biases, fairness, and stability. Our findings highlight the potential and challenges of integrating LLMs into recommender systems, paving the way for future research and improvements in this domain.
@@
The structure of the paper is organized as follows. Section 2 presents the related work, focusing on the exploration of fairness in recommendation systems (RS) and pre-trained language models (LMs), as well as their applications in enhancing RS. Section 3 is dedicated to presenting a thorough evaluation of ChatGPT. This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1.1), the repeated experiments for stability analysis (Section 3.1.2), an understanding of the system role in ChatGPT-based RecLLM (Section 3.1.3), and the promotion of fairness (Section 3.1.4). Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.1.5). Following this, Section 4 outlines the experimental setup employed in our study. The results and key findings of these experiments are thoroughly discussed in Section 5. The paper concludes with Section 6, where we summarize our findings and outline potential avenues for future research.
@@ 2.1. Fairness in Recommender Systems
Fairness has become a pivotal topic in AI, gaining scrutiny across branches of trustworthy AI, including security, privacy, and explainability. Fairness in recommender systems has gained significant attention due to their multi-stakeholder nature (Deldjoo et al., 2023; Ekstrand et al., 2019). Unfairness, even in its minimal form, can adversely impact various stakeholders, including consumers, producers, system designers, supply chains, and even the environment – the latter are often referred to as ‘side-stakeholder’.
@@
The body of literature on fairness in RS is diverse, covering multiple perspectives and dimensions. Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), categorize these aspects into several dimensions, either orthogonal or partially orthogonal. Key dimensions recognized in the literature include the main stakeholder in question (e.g., consumer vs. producer), the target benefits associated with each (such as effectiveness vs. item exposure), the granularity of sensitive groups for assessing fairness (individual vs. group level), and other dimensions including the core definition of fairness, temporal aspects, and more. These dimensions offer a comprehensive view of the fairness landscape in RS, as detailed in these surveys.
@@
To position this study within the literature of fair recommender systems, we introduced Table 1. This table is designed to categorize existing literature along two principal dimensions: the stakeholder in question (consumer vs. producer) and the nature of the core RS under scrutiny (traditional RS vs. those based on recent advancements in LLMs). We briefly review these dimensions in the following.
@@
Core RS under scrutiny. We distinguish between ‘traditional’ RS and those enhanced by ‘RecLLM’. This distinction is crucial because, while RecLLM, such as those based on GPT-like architectures, promise to significantly advance RS landscape with more nuanced and personalized recommendations, they also raise concerns about inherent biases. The extensive and unregulated nature of internet data used to train LLMs raises concerns about biases against specific races, genders, and other sensitive attributes. For example, if an LLM is trained on e-commerce data where men’s products are described more positively than women’s products, it might skew recommendations towards male-oriented items. Therefore, measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies. We can briefly review the research on fairness in recommendation systems according to Table 1 as follows:
@@
  • Traditional RS. This column lists studies that have focused on traditional methods of recommender systems (Deldjoo et al., 2018). They primarily rely on CF algorithms (possibly using side information of users and items) without the advanced natural language processing capabilities of LLMs. Within these works, some are dedicated to building evaluation frameworks for evaluating RS unfairness, while others focus on developing various mitigation strategies.
@@
  • RecLLM: This column spotlights the burgeoning research domain that merges language models (LMs) with RS representing a significant shift towards harnessing advanced NLP techniques to enhance the accuracy and relevance of recommendations. These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023; Zhang et al., 2023a). Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023; Zhang et al., 2023a), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
@@
Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023a; Deldjoo and Di Noia, 2024; Deldjoo, 2024)
@@
Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023)
@@
Stakeholder. As mentioned earlier, a major aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides (Deldjoo et al., 2022; Rahmani et al., 2024; Naghiaei et al., 2022). Within each of these segments, as you can observe, we can further categorize the literature based on which sensitive attribute groups are defined. For example, sensitive attributes such as the demographics of consumers (age, gender) and producers, as well as the popularity of items (on the produce side) are quite common focal points.
  • Consumer Fairness. This category is subdivided into various attributes like Activity, Demographics, Merits, and Others. It includes studies that focus on ensuring fairness among consumers of the recommender system based on these attributes. For instance, ensuring that recommendations are not biased towards a particular demographic consumer group.
@@
The study by Zhang et al. (2023a) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness. Their work introduces a novel benchmark, FaiRLLM, for evaluating the fairness of RecLLM, highlighting ChatGPT’s biases towards certain sensitive user attributes in music and movie recommendations. The work by Deldjoo et al. (Deldjoo and Di Noia, 2024; Deldjoo, 2024) address the shortcomings by scrutinizing whether changes in recommendations, due to the inclusion of sensitive attributes, result in unfairness. It also explores better ways to normalize and improve the fairness of these models. Conversely, the work by Li et al. (2023) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation. This study investigates performance of ChatGPT in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection. However, these studies do not address the intricacies of prompt engineering in RecLLMs, nor do they comprehensively address the various forms of biases studies in the current study. Our research goes beyond examining provider fairness based on item unfairness and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.